find three consecutive integers such that the sum of twice the second and three time the third is five less than six times the first
Answer by lwsshak3(11628)

find three consecutive integers such that the sum of twice the second and three time the third is five less than six times the first.
***
let x=1st consecutive integer
let x+1=2nd consecutive integer
let x+2=3rd consecutive integer
..
2(x+1)+3(x+2)=6x-5
2x+2+3x+6=6x-5
x=13
x+1=14
x+2=15xx
1st consecutive integer=13
2nd consecutive integer=14
3rd consecutive integer=15